OpenStudy (zolock):

Show all your work. 1. Solve each equation, if possible. Write irrational numbers in simplest radical form. Describe the strategy you used to get your solution and tell why you chose that strategy. a. x^2 +4 =0 b. x^2 -6x +1=0

OpenStudy (zolock):

what happened was it wrong or something

OpenStudy (anonymous):

a. x^2 + 4 = 0 x^2 = -4 x = ± √-4 x = ± √(-1 * 4) x = ± 2i b. x^2 - 6x + 1 = 0 d = (-6)^2 - 4(1)(1) = 32 x = (6 ± √32) / 2 x = 3 ± 2√2

OpenStudy (anonymous):

sorry i look at the problem way wrong, I thought your question had 49!!! sorry haha

OpenStudy (anonymous):

thats right^^^

OpenStudy (zolock):

it`s ok thanks

OpenStudy (anonymous):

your welcome.

OpenStudy (zolock):

Can u please Describe the strategy you used to get your solution and tell why you chose that strategy.